The electric potential at the surface of a charged sphere is 10 V at its surface. Find the value of potential and electric field at its centre
1
Expert's answer
2021-07-19T15:28:34-0400

Gives

"V_s=\frac{kQ}{r}"

Q charge distributed by surface

"V_s=10V"

Centr potential

"V_c=\frac{kQ}{r}" Where r<R


Then "V_c=10V"

Electric field

"E_s=\frac{Kq}{R^2}" Where R=r

"E_{in}=\frac{kQr}{R^3}"

Centre electric field

Put value

r=0

"E_c=0"